OpenStudy (anonymous):

is the mixed number 2 2/3 in simplest form?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i say yes it is in simplest form

OpenStudy (anonymous):

In the form 8/3 it may use up less ink when printed.

OpenStudy (anonymous):

is 8 2/5 - 6 3/5, does any fraction need to be renamed?

OpenStudy (anonymous):

8 2/5 - 6 3/5 = 8+2/5 - 6+3/5 = 3 I guess the answer is 3 is a "whole number" or "integer"

OpenStudy (anonymous):

Sorry. Ignore my last posting.

OpenStudy (anonymous):

\[(8 +2/5)-(6 +3/5)=\frac{42}{5}-\frac{33}{5}=\frac{9}{5} \]Normally I would leave an answer like 9/5, "nine fifths", the way it is.

OpenStudy (anonymous):

is 1 4/5 fraction in simplest form

OpenStudy (anonymous):

As a mixed number I would say yes. For instance, 1 8/10 can be reduced to 1 4/5 and no further. 4 and 5 have no common divisor.